Cactus Cookies incurred the following costs this month to manufacture 1,000 jumbo chocolate chip cookies Chocolate chips Flour Butter Sugar Other ingredients Labor wages Rent on manufacturing space Utilities (paid at a fixed monthly rate) Depreciation on kitchen equipment 58.00 $30.00 $61.50 $30.00 $35.00 $ 300.00 $ 300.00 54.00 $ 82.00

d. The fixed cost per cooksie this month. Round your answer to 2 decimal places ) e. The total variable cost if Cactus Cookies had manufactured 2,000 cookies this month. (Do not round your intermediate calculations.) f. The total fixed cost if Cactus Cookies had manufactured 2,000 cookies this month 9. The variable cost per cookie i Cactus Cookies had manufactured 2 00 cokis ie month (Round your answer to 4 decimal places)

Need help with D, E, F, G
Answer #1
d. Fixed Cost per cookie this month:
Fixed Cost items Amount $
Rent on manufacturing space 300
Utilities (paid at a fixed monthly rate) 54
Depreciation on kitchen equipment 82
Total Fixed Costs 436
Fixed cost per cookie = Total fixed costs/Cookies produced
Fixed cost per cookie = 436/1000 = $0.436 or $0.44 per cookie
e. Total variable cost if 2000 cookies manufactured this month:
Variable Cost items Amount $
Chocolate chips 58
Flour 30
Butter 61.5
Sugar 30 0.5145
Other ingredients 35
Labor wages 300
Total Variable cost to manuf. 1000 cookies 514.5
Total Variable cost to manuf. 2000 cookies= TVC to manuf. 1000 cookies *2
Total Variable cost to manuf. 2000 cookies=514.50 * 2 = $1029
f. Total fixed cost, had manuf. 2000 cookies = $436 (assumption not operating on capacity)
g. Variable cost per cookie if manuf. 2000 this month = $1029 / 2000 = $ 0.5145 or $0.52 per cookie